A 2.50 kg mass is pushed 10.0 m up a 30° incline. The work done is 125 J. What force is applied parallel to the incline to move the object?

a. 10.8 N

b. 12.5 N

c. 14.4 N

d. 24.5 N
Answer #1

Answer : --

  • let this force = F
  • Now the Object is acting by following force
  • mgcos\theta downward
  • mgsin\theta parallel to the incline but opposite to F
  • so the total parallel force = F- mgsin\theta
  • the work ddone by this force is given by
  • W = (F- mgsin\theta )d
  • 125= (F-2.5 \times 9.8 \times sin(30) )10
  • 12.5=F - 2.5 \times 9.8 \times sin(30)
  • F= 12.5+2.5 \times 9.8 \times \frac{1}{2}
  • F= 24.75N
  • so the option no. (4) is correct.